OpenStudy (anonymous):

Solve by Substitution method. 10x+3y=8 2x + y=2 Step by step also please.

OpenStudy (anonymous):

same here. 10x+3y=8 y=2-2x. Substitute (2-2x) for Y in the 10x+3y=8 10x+3(2-2x)=8 and solve for x

OpenStudy (anonymous):

from the second equation 2x+y = 2 y = 2-2x then substitute y to the first equation 10x+3y=8 10x+3(2-2x) = 8 10x+6-6x=8 10x-6x=8-6 4x=2 x=0.5

OpenStudy (anonymous):

then plug x into one of those equations to get y if you plug x to the first equation: 10x+3y=8 10(0.5) +3y=8 5+3y=8 3y=3 y=1 or if you want to put x to the second equation: 2x + y=2 2(0.5) +y=2 1+y = 2 y=1 you'll get the same answer

OpenStudy (anonymous):

I have the answer x=1/2 y= 1

OpenStudy (anonymous):

that's right